OpenStudy (anonymous):

/Users/admin/Desktop/Screen Shot 2014-12-05 at 09.55.27 am.png

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy :) sorry but that file can only be viewed from your personal computer to let other people see it, click the "attach file" button below :)

OpenStudy (sleepyhead314):

well the lines are divided into 10 sections so that means that each small line is worth 1/10 or 0.1 so 0.3 = 3 lines from 1 7/5 = 14/10 = 4 lines from 1 1 4/5 = 1 8/10 = 8 lines from 1 do you understand that? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so A

OpenStudy (sleepyhead314):

yes :) that is correct! :D

OpenStudy (anonymous):

thank so much!!

OpenStudy (sleepyhead314):

you're welcome ^_^